# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ 8.5 Minimality by anyon count/quantum dimension
As we emphasized, the observation of a fractional $\sigma_{H}$ implies the existence of a set of Abelian anyons. We now pose the question of what the “minimal” topological order is in a system with a given fractional $\sigma_{H}$. As some of the authors argued in a recent paper [30] two reasonable definitions are that minimal theories have the smallest: 1) number of anyons, $N_{{\cal T}}$, or 2) total quantum dimension, $\mathcal{D}_{{\cal T}}$, among all theories consistent with the IR symmetries and anomalies. We will determine minimal theories for FQH systems through either definition below.
@@
a Single layer spinc, even $q$
@@
We start with $N_{{\cal T}}$. Due to charge conservation for any anyon $x\in\mathcal{T}$, $x\times v^{k}$ for $k=1,\cdots,q-1$ are all distinct from each other because $v^{k}$ carries charge $pk/q$. This implies that $N_{\mathcal{T}}$ must be a multiple of $q$. Together with the bounds found above, we conclude $N_{\mathcal{T}}=3q$.
We next group anyons in $\mathcal{T}$ according to their braiding with $v^{q}$. We take $\mathcal{T}_{0}$ ($\mathcal{T}_{1}$) to be those with braiding phase 00 ($1/2$). Clearly ${\cal V}^{2q,2p}\subset\mathcal{T}_{0}$. Also, since $\mathcal{T}_{1}$ is non-empty, the argument above implies there are at least $q$ anyons in $\mathcal{T}_{1}$. Since $N_{\mathcal{T}}=3q$ this tells us that these inequalities are equalities and $\mathcal{T}_{0}={\cal V}^{2q,2p}$, $\mathcal{T}_{1}=\{x\times v^{k}|k=0,1,\dots,q-1\}$. In particular, this implies $x\times v^{q}c=x$ [^37] Here we have not identified $c\sim 1$ to make it manifest that this fusion rule obeys the spin/charge condition. It is easy to see that $\mathcal{T}=\mathcal{T}_{0}\oplus\mathcal{T}_{1}$ gives a $\mathbb{Z}_{2}$ grading on $\mathcal{T}$, using Lemma D.2, we see that $2q={\cal D}^{2}_{{\cal T}_{0}}={\cal D}_{{\cal T}_{1}}^{2}=qd_{x}^{2}$ and thus that $d_{x}=\sqrt{2}$. The total quantum dimension squared is ${\cal D}^{2}_{{\cal T}}=2{\cal D}^{2}_{{\cal T}_{0}}=4q$.
To construct this theory, we tentatively violate the spinc condition and gauge the one-form symmetry generated by the charge one boson $v^{q}$. Then, $\mathcal{T}$ is reduced to ${\cal V}^{q,p}$. To recover the parent TQFTs, we gauge the dual zero-form $\mathbb{Z}_{2}^{(0)}$ symmetry acting in the child TQFT ${\cal A}^{q,p}$. The action on the anyons is determined such that after the gauging, ${\cal A}^{q,p}$ becomes ${\cal V}^{2q,2p}$. One choice for $\mathcal{T}$ is ${\cal A}^{4q,p}$ with vison $v=a_{2}$. All other choices of $\mathcal{T}$ can be obtained by tensoring invertible theories when gauging. In this case, the invertible theories are fermionic SPT phases with global $\mathbb{Z}_{2}^{(0)}$ symmetry [125]. Gauging the $\mathbb{Z}_{2}^{(0)}$ symmetry in the fermionic SPT phases gives ${\rm Spin}(n)_{1}$, with $\psi c$ being the gauge charge. We thus obtain the full list of candidates for $\mathcal{T}$, by tensoring ${\cal A}^{4q,p}$ and ${\rm Spin}(n)_{1}$, and then gauging the diagonal $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $\psi a_{2q}c$:
@@
We stop to note that all these minimal TQFTs can also be understood through the parton construction. Write the electron as $c=fb$, where $f$ is the neutral fermion and $b$ is a charge-1 boson. The $b$ bosons are at filling $p/q$, and the unique minimal TQFT for them is $\mathcal{V}^{q,p}$. [^38] We show this in section 8.6, when discussing minimal orders of bosonic FQH states. We then let the neutral fermions form a chiral topological superconductor with Chern number $n$.
@@
The problem then reduces to finding which of these can be consistent with (anomalous) time reversal. Let $\sigma$ be one of the two anyons in $\mathsf{Pf}^{f}_{2,1,n}$ that have charge $Q(\sigma)=1/4$. It can be checked that $h(\sigma)=(n+1)/16$. Then since the full symmetry is $\mathrm{U}(1)\rtimes\mathbb{Z}_{2}^{\cal T}$, i.e., time reversal preserves the $\mathrm{U}(1)$ charge, ${\cal T}\sigma=\sigma$ or $v^{2}c\times\sigma$, with the latter being distinct from $\sigma$ for even $n$. Time reversal is anti-unitary, so $h({\cal T}\sigma)=-h(\sigma)$. This, along with ${\cal T}\sigma=\sigma$ or $v^{2}c\times\sigma$, means that $2h(\sigma)=0\pmod{1}$. Since $h(\sigma)=(n+1)/16$, the only time-reversal-invariant TQFT among this list is the T-Pfaffian TQFT, corresponding to $\mathsf{Pf}^{f}_{2,1,-1}\simeq\mathsf{Pf}^{f}_{2,1,7}$ in the above notation. Moreover, this action of time reversal is anomalous, as can be seen from the fact that the vison $v$ carries charge $1/2$; in a non-anomalous time-reversal-invariant 2+1d system, the vison must be charge-neutral because $\sigma_{H}=0$.
@@
The previous discussion can be generalized to multi-layer FQH systems. Here “layer” should be understood generally as an independent, discrete quantum number. For example, it can also be the spin, or the valley degree of freedom in graphene or moiré bands. The key assumption is that each layer comes with a separate $\mathrm{U}(1)$ charge, which are all preserved in the ground state. Note that this requirement means that we will not allow “interlayer” tunneling. If we did, it would break the two $\mathrm{U}(1)$ symmetries to a single $\mathrm{U}(1)$, taking us back to our single-layer discussion.
@@
In the bilayer example, the spin/charge relation requires all local anyons to be bosons (fermions) if they have even (odd) total charge, $Q_{1}+Q_{2}$. Since $v_{1}^{2},v_{2}^{2}$ are both bosons with total charge $1$, the spin/charge relation means $v_{1}^{2},v_{2}^{2}\neq 1$. Thus $\ell_{1}$ and $\ell_{2}$ must be even. There must then be at least $8$ anyons in $\mathcal{S}\subseteq\mathcal{T}$, lower bounding the anyon count and quantum dimension squared. [^40] The lower bound is $8=4\times 4/2$ and not $16$ because of the possibility that $v_{1}^{2}=v_{2}^{2}$, as seen in the example of $D(\mathbb{Z}_{4})$. On the other hand, there is a known candidate TQFT
@@
This upper bound immediately tells us that in any minimal theory, we must have $\ell_{1}=\ell_{2}=2$ and $v_{1}^{2}v_{2}^{2}=1$. Just as in the single-layer case, we can gauge the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $v_{1}^{2}=v_{2}^{2}$ to produce the child TQFT $\mathcal{A}^{2,1}\boxtimes\mathcal{A}^{2,1}$. We then gauge the dual $\mathbb{Z}_{2}^{(0)}$ 00-form symmetry in this child TQFT, with the ambiguity of stacking fermionic SPT states protected by $\mathbb{Z}_{2}^{(0)}\times\mathbb{Z}_{2}^{(0),f}$. As in the single-layer case, this gives us the possible parent TQFTs
@@
8.6 Single layer bosonic theories
@@
However, when both $q$ and $p$ are odd, ${\cal V}^{q,p}$ is not a valid bosonic theory. Thus $\ell$ must be even. Therefore, the minimal theory has $\ell=2$ and includes [^41] In the first step, we used the fact that for odd $q$, as groups $\mathbb{Z}_{2q}=\mathbb{Z}_{2}\otimes\mathbb{Z}_{q}$. And then, there cannot be a mixed anomaly between the two factors. In the second step, we used the relations derived in Sec. 7 and the fact that $p$ and $q$ are odd. It should also be noted that the vison in the left-hand side is the product of the generators of the two factors in the right-hand side. $\mathcal{V}^{2q,2p}\cong\mathcal{A}^{2,2pq}\boxtimes\mathcal{A}^{q,4p}\cong\mathcal{A}^{2,2}\boxtimes\mathcal{A}^{q,4p}$. Since this is not a complete theory, the full theory does not factorize, and there must be additional lines charged under the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry. As an example, for $p=q=1$ (corresponding to the bosonic $\nu=1$ state) the minimal symmetry is ${\cal A}^{2,2}$. It includes the identity and a nontrivial vison with spin $1/2$. This is not a complete TQFT. It can be completed in various ways including Spin$(n)_{1}$ and ${\rm SU}(2)_{4k+2}$. This leads to infinitely many endpoint theories, similar to the single-layer spinc case with even $q$.
@@ 9 Summary and discussion
Before stating our results in the general case, let us specialize to the question of minimal topological order in electronic systems where the fractional quantum Hall effect has been observed in experiments. First, note that $(1,v,v^{2},.....v^{q-1})$ are all distinct quasiparticles. Second, $v^{q}$ braids trivially with $v^{j}$ for all $j$. For odd-denominator states (i.e., odd $q$), it has charge $p$ and Fermi (Bose) statistics when $p$ is odd (even). The minimal topological order is one where $v^{q}$ is simply identified with a local particle. Thus, the minimal topological order is unique and has exactly $q$ anyons (and is denoted ${\cal V}^{q,p}$). For even-denominator states, $v^{q}$ is an odd-charged boson, which therefore cannot be local in an electronic system. There must therefore be anyons distinct from those generated by the vison with which $v^{q}$ braids non-trivially. We showed that in the minimal topological order for even $q$, $v^{2q}$ can be identified with a local particle. Though the minimal order in this case is not unique, there are a small number of them that are all Pfaffian states. In particular, the minimal orders by anyon count have non-Abelian anyons with which $v^{q}$ braids non-trivially.
Our conclusions about the minimal theories for a given $\sigma_{H}$ (by anyon count or quantum dimension) are summarized in Table 2.
Spinc, $\sigma_{H}=p/q$
$q$ odd $\mathcal{V}^{q,p}$ $\mathcal{V}^{q,p}$ $q$ $q$
$q$ even $\mathcal{V}^{2q,2p}$ $\mathsf{Pf}^{f}_{q,p,n}$ $4q$ $4q$ ($3q$) for $n$ even (odd)
Bosons, $\sigma_{H}=p/q$
$pq$ even $\mathcal{V}^{q,p}$ $\mathcal{V}^{q,p}$ $q$ $q$
$pq$ odd $\mathcal{V}^{2q,2p}=\mathcal{A}^{2,2}\boxtimes\mathcal{A}^{q,p+q}$ $\mathsf{Pf}^{b}_{q,p,n}$ $4q$ $4q$ ($3q$) for $n$ even (odd)
Table 2: A summary of our results. This table gives the list of possible minimal TQFTs by quantum dimension $\mathcal{D}^{2}_{{\cal T}}$ and anyon count $N_{{\cal T}}$ for each FQH setup. See App. A for a description of the notation. In all the examples, we find that the minimal TQFTs by anyon count are a subset of the minimal TQFTs by quantum dimension; if there are non-Abelian options, as in the case of the odd $n$ theories above, they are precisely these options.
@@
    If all the fundamental degrees of freedom are bosons, then our spacetime does not have to be a spin manifold. And if it is a spin manifold, there is no dependence on its spin structure. In that case, shifting $\sigma_{H}$ by an even integer corresponds to adding a local counter-term and does not affect the bulk dynamics. Even though it can affect the boundary dynamics, we will ignore it and identify
@@
    If the fundamental degrees of freedom include also fermions, the theory can be formulated only on spin manifolds, and the physical answers can depend on the choice of spin structure. In that case, shifting $\sigma_{H}$ by an arbitrary integer corresponds to adding a local counter-term and does not affect the bulk dynamics. Even though it can affect the boundary dynamics, we will ignore it and identify
@@
Thus, we see that the vast majority of observed fractional quantum Hall states are described by minimal TQFTs. However, based on numerical calculations, a few non-minimal TQFTs have been proposed as well, primarily in the second Landau level. The most prominent of such states is one at Landau level filling $\nu=\frac{12}{5}$. Numerical studies [116, 117, 118, 119, 120] (for the Coulomb interaction) appear consistent with the particle-hole conjugate $\bar{R}_{1,3}$ of the Read-Rezayi state $R_{1,3}$, though a Jain state ${\cal V}^{5,2}$ is allowed at that filling. As $\bar{R}_{1,3}=\bar{{\cal F}_{3}}\boxtimes\mathcal{V}^{5,2}$, the Read-Rezayi state has 5 more anyons than what is needed for minimality. In experiments in GaAs [149, 150], the corresponding state at $\nu=\frac{13}{5}$ is not observed, an effect that has been attributed to Landau level mixing [119, 120]. For a few other proposed non-minimal states in the second Landau level, see [151, 152, 153] though these have not been studied as extensively as the possibility of the $\bar{R}_{1,3}$ state at filling $\frac{12}{5}$.
Interesting possible examples of non-Jain states occur at fillings such as $\frac{8}{17}$ or $\frac{7}{13}$ [154]. Though conventional Jain states are allowed, these have been proposed to be so-called “daughter states” of parent quantum Hall states at the proximate filling $\nu=\frac{1}{2}$ [155]. The parent $\nu=\frac{1}{2}$ states are Pfaffians: at this filling, the composite fermions interact with a zero effective magnetic field, and the Pfaffian states result when Cooper pairs of composite fermions are condensed. When the filling is moved away from $1/2$, the composite fermion Cooper pairs interact with a non-zero effective magnetic field, which suppresses their condensation. Then, the daughter states appear at fillings for which the composite fermion Cooper pairs form bosonic integer quantum Hall states rather than condensates [156]. Different Pfaffian parent states have daughters at different fillings, and thus the observation of these daughters has been suggested to be a fingerprint of which Pfaffian state is realized at half-filling [156, 157], a question that so far has not yet been settled. From the point of view in the present paper, these daughter states are minimal at their fillings. This is because they can all be described as Jain states tensored with invertible electrically neutral $E_{8}$ states [157]. Thus, the difference with the Jain states is not at the level of anyon content, and these daughter states have the minimal number of anyons at their fillings.

### v3

@@ 8.5 Minimality by anyon count/quantum dimension
As we emphasized, the observation of a fractional $\sigma_{H}$ implies the existence of a set of Abelian anyons. We now pose the question of what the “minimal” topological order is in a system with a given fractional $\sigma_{H}$. As some of the authors argued in a recent paper [30] two reasonable definitions are that minimal theories have the smallest: 1) number of anyons $N_{{\cal T}}$ or 2) total quantum dimension $\mathcal{D}_{{\cal T}}$, among all theories consistent with the IR symmetries and anomalies. We will determine minimal theories for FQH systems through either definition below.
@@
a Single-layer spinc, even $q$
@@
We start with $N_{{\cal T}}$. Because of charge conservation for any anyon $x\in\mathcal{T}$, $x\times v^{k}$ for $k=1,\cdots,q-1$ are all distinct from each other because $v^{k}$ carries charge $pk/q$. This implies that $N_{\mathcal{T}}$ must be a multiple of $q$. Together with the bounds found above, we conclude $N_{\mathcal{T}}=3q$.
We next group anyons in $\mathcal{T}$ according to their braiding with $v^{q}$. We take $\mathcal{T}_{0}$ ($\mathcal{T}_{1}$) to be those with braiding phase 0 ($1/2$). Clearly ${\cal V}^{2q,2p}\subset\mathcal{T}_{0}$. Also, since $\mathcal{T}_{1}$ is non-empty, the argument above implies there are at least $q$ anyons in $\mathcal{T}_{1}$. Since $N_{\mathcal{T}}=3q$ this tells us that these inequalities are equalities and $\mathcal{T}_{0}={\cal V}^{2q,2p}$, $\mathcal{T}_{1}=\{x\times v^{k}|k=0,1,\dots,q-1\}$. In particular, this implies $x\times v^{q}c=x$. [^37] Here we have not identified $c\sim 1$ to make it manifest that this fusion rule obeys the spin/charge condition. It is easy to see that $\mathcal{T}=\mathcal{T}_{0}\oplus\mathcal{T}_{1}$ gives a $\mathbb{Z}_{2}$ grading on $\mathcal{T}$, using Lemma D.2, we see that $2q={\cal D}^{2}_{{\cal T}_{0}}={\cal D}_{{\cal T}_{1}}^{2}=qd_{x}^{2}$ and thus that $d_{x}=\sqrt{2}$. The total quantum dimension squared is ${\cal D}^{2}_{{\cal T}}=2{\cal D}^{2}_{{\cal T}_{0}}=4q$.
To construct this theory, we tentatively violate the spinc condition, as in Fig. 4, and gauge the one-form symmetry generated by the charge one boson $v^{q}$. Then, $\mathcal{T}$ is reduced to ${\cal V}^{q,p}$. To recover the parent TQFTs, we gauge the dual zero-form $\mathbb{Z}_{2}^{(0)}$ symmetry acting in the child TQFT ${\cal A}^{q,p}$. The action on the anyons is determined such that after the gauging, ${\cal A}^{q,p}$ becomes ${\cal V}^{2q,2p}$. One choice for $\mathcal{T}$ is ${\cal A}^{4q,p}$ with vison $v=a_{2}$. All other choices of $\mathcal{T}$ can be obtained by tensoring invertible theories when gauging. In this case, the invertible theories are fermionic SPT phases with global $\mathbb{Z}_{2}^{(0)}$ symmetry [123]. Gauging the $\mathbb{Z}_{2}^{(0)}$ symmetry in the fermionic SPT phases gives ${\rm Spin}(n)_{1}$, with $\psi c$ being the gauge charge. We thus obtain the full list of candidates for $\mathcal{T}$, by tensoring ${\cal A}^{4q,p}$ and ${\rm Spin}(n)_{1}$, and then gauging the diagonal $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $\psi a_{2q}c$,
@@
We stop to note that all these minimal TQFTs can also be understood through the parton construction. Write the electron as $c=fb$, where $f$ is the neutral fermion and $b$ is a charge-1 boson. The $b$ bosons are at filling $p/q$, and the unique minimal TQFT for them is $\mathcal{V}^{q,p}$. [^38] We show this in Sec. 8.6, when discussing minimal orders of bosonic FQH states. We then let the neutral fermions form a chiral topological superconductor with Chern number $n$.
@@
The problem then reduces to finding which of these can be consistent with (anomalous) time reversal. Let $\sigma$ be one of the two anyons in $\mathsf{Pf}^{f}_{2,1,n}$ that have charge $Q(\sigma)=1/4$. It can be checked that $h(\sigma)=(n+1)/16$. Then since the full symmetry is $\mathrm{U}(1)\rtimes\mathbb{Z}_{2}^{\cal T}$, i.e., time reversal preserves the $\mathrm{U}(1)$ charge, ${\cal T}\sigma=\sigma$ or $v^{2}c\times\sigma$, with the latter being distinct from $\sigma$ for even $n$. Time reversal is anti-unitary, so $h({\cal T}\sigma)=-h(\sigma)$. This, along with ${\cal T}\sigma=\sigma$ or $v^{2}c\times\sigma$, means that $2h(\sigma)=0\pmod{1}$. Since $h(\sigma)=(n+1)/16$, the only time-reversal-invariant TQFT among this list is the T-Pfaffian TQFT, corresponding to $\mathsf{Pf}^{f}_{2,1,-1}\simeq\mathsf{Pf}^{f}_{2,1,7}$ in the above notation. Moreover, this action of time reversal is anomalous, as can be seen from the fact that the vison $v$ carries charge $1/2$; in a non-anomalous time-reversal-invariant 2+1d system, the vison must be charge neutral because $\sigma_{H}=0$.
@@
The previous discussion can be generalized to multilayer FQH systems. Here “layer” should be understood generally as an independent, discrete quantum number. For example, it can also be the spin, or the valley degree of freedom in graphene or moiré bands. The key assumption is that each layer comes with a separate $\mathrm{U}(1)$ charge, which are all preserved in the ground state. Note that this requirement means that we will not allow “interlayer” tunneling. If we did, it would break the two $\mathrm{U}(1)$ symmetries to a single $\mathrm{U}(1)$, taking us back to our single-layer discussion.
@@
In the bilayer example, the spin/charge relation requires all local anyons to be bosons (fermions) if they have even (odd) total charge, $Q_{1}+Q_{2}$. Since $v_{1}^{2},v_{2}^{2}$ are both bosons with total charge $1$, the spin/charge relation means $v_{1}^{2},v_{2}^{2}\neq 1$. Thus $\ell_{1}$ and $\ell_{2}$ must be even. There must then be at least eight anyons in $\mathcal{S}\subseteq\mathcal{T}$, lower bounding the anyon count and quantum dimension squared. [^40] The lower bound is $8=4\times 4/2$ and not $16$ because of the possibility that $v_{1}^{2}=v_{2}^{2}$, as seen in the example of $D(\mathbb{Z}_{4})$. On the other hand, there is a known candidate TQFT
@@
This upper bound immediately tells us that in any minimal theory, we must have $\ell_{1}=\ell_{2}=2$ and $v_{1}^{2}v_{2}^{2}=1$. Just as in the single-layer case, we can gauge the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $v_{1}^{2}=v_{2}^{2}$ to produce the child TQFT $\mathcal{A}^{2,1}\boxtimes\mathcal{A}^{2,1}$. We then gauge the dual $\mathbb{Z}_{2}^{(0)}$ 0-form symmetry in this child TQFT, with the ambiguity of stacking fermionic SPT states protected by $\mathbb{Z}_{2}^{(0)}\times\mathbb{Z}_{2}^{(0),f}$. As in the single-layer case, this gives us the possible parent TQFTs
@@
8.6 Single-layer bosonic theories
@@
[…]
@@ 9 Summary and discussion
Before stating our results in the general case, let us specialize to the question of minimal topological order in electronic systems where the fractional quantum Hall effect has been observed in experiments. First, note that $(1,v,v^{2},.....v^{q-1})$ are all distinct quasiparticles. Second, $v^{q}$ braids trivially with $v^{j}$ for all $j$. For odd-denominator states (i.e., odd $q$), it has charge $p$ and Fermi (Bose) statistics when $p$ is odd (even). The minimal topological order is one where $v^{q}$ is simply identified with a local particle. Thus, the minimal topological order is unique and has exactly $q$ anyons (and is denoted ${\cal V}^{q,p}$). For even-denominator states, $v^{q}$ is an odd-charged boson, which therefore cannot be local in an electronic system. There must therefore be anyons distinct from those generated by the vison with which $v^{q}$ braids nontrivially. We showed that in the minimal topological order for even $q$, $v^{2q}$ can be identified with a local particle. Though the minimal order in this case is not unique, there are a small number of them that are all Pfaffian states. In particular, the minimal orders by anyon count have non-Abelian anyons with which $v^{q}$ braids nontrivially.
Our conclusions about the minimal theories for a given $\sigma_{H}$ (by anyon count or quantum dimension) were summarized in Table 2 in our summary of results.
@@
    If all the fundamental degrees of freedom are bosons, then our spacetime does not have to be a spin manifold. And if it is a spin manifold, there is no dependence on its spin structure. In that case, shifting $\sigma_{H}$ by an even integer corresponds to adding a local counterterm and does not affect the bulk dynamics. Even though it can affect the boundary dynamics, we will ignore it and identify
@@
    If the fundamental degrees of freedom include also fermions, the theory can be formulated only on spin manifolds, and the physical answers can depend on the choice of spin structure. In that case, shifting $\sigma_{H}$ by an arbitrary integer corresponds to adding a local counterterm and does not affect the bulk dynamics. Even though it can affect the boundary dynamics, we will ignore it and identify
@@
Thus, we see that the vast majority of observed fractional quantum Hall states are described by minimal TQFTs. However, based on numerical calculations, a few non-minimal TQFTs have been proposed as well, primarily in the second Landau level. The most prominent of such states is one at Landau level filling $\nu=\frac{12}{5}$. Numerical studies [113, 115, 116, 117, 118] (for the Coulomb interaction) appear consistent with the particle-hole conjugate $\bar{R}_{1,3}$ of the Read-Rezayi state $R_{1,3}$, though a Jain state ${\cal V}^{5,2}$ is allowed at that filling. As $\bar{R}_{1,3}=\bar{{\cal F}_{3}}\boxtimes\mathcal{V}^{5,2}$, the Read-Rezayi state has five more anyons than what is needed for minimality. In experiments in GaAs [147, 148], the corresponding state at $\nu=\frac{13}{5}$ is not observed, an effect that has been attributed to Landau level mixing [117, 118]. For a few other proposed nonminimal states in the second Landau level, see [149, 150, 151] though these have not been studied as extensively as the possibility of the $\bar{R}_{1,3}$ state at filling $\frac{12}{5}$.
Interesting possible examples of non-Jain states occur at fillings such as $\frac{8}{17}$ or $\frac{7}{13}$ [152]. Though conventional Jain states are allowed, these have been proposed to be so-called “daughter states” of parent quantum Hall states at the proximate filling $\nu=\frac{1}{2}$ [153]. The parent $\nu=\frac{1}{2}$ states are Pfaffians: at this filling, the composite fermions interact with a zero effective magnetic field, and the Pfaffian states result when Cooper pairs of composite fermions are condensed. When the filling is moved away from $1/2$, the composite fermion Cooper pairs interact with a non-zero effective magnetic field, which suppresses their condensation. Then, the daughter states appear at fillings for which the composite fermion Cooper pairs form bosonic integer quantum Hall states rather than condensates [154]. Different Pfaffian parent states have daughters at different fillings, and thus the observation of these daughters has been suggested to be a fingerprint of which Pfaffian state is realized at half-filling [154, 155], a question that so far has not yet been settled. From the point of view in the present paper, these daughter states are minimal at their fillings. This is because they can all be described as Jain states tensored with invertible electrically neutral $E_{8}$ states [155, 34]. Thus, the difference with the Jain states is not at the level of anyon content, and these daughter states have the minimal number of anyons at their fillings.